OpenStudy (anonymous):

Quick graphing question, thanks for helping! I posted the graph in the comments. I need to write the a coordinate in polar form, but make the radius a positive value. I also need to write point B in polar form as well. Thank You, this topic is really confusing. :P

OpenStudy (anonymous):

@mathstudent55 @Nnesha @thomaster @sammixboo @pooja195 @freckles @inkyvoyd @k_lynn @mathmath333 @ybarrap @AkashdeepDeb @tanya123

OpenStudy (anonymous):

yup

OpenStudy (ybarrap):

Using Euler's Identity: $$ \large Ae^{i\theta}=A(\cos \theta +i \sin \theta ) $$  Assuming each ring is 1 unit. Does this make sense?